1. What is the pH of a 0.350 M solution of lithium citrate? (Ka Citric Acid = 7.4 x 10-4

2.Consider the following reaction:

PCl3 (g)+ Cl2 (g)   PCl5 (g)                                      

K = 4.30 x 10-6

Initially, 2.50 M PCl3 and 1.40 M Cl2 are placed into a container and allowed to reach equilibrium. Determine the equilibrium concentrations of all the species


1
Expert's answer
2021-02-02T01:25:39-0500

1. PH = 8.33


2 . PCl3 + Cl2 "\to" PCl5

I = 2.50...1.40 .......0

C = -x ..... -x ........+x

E =(2.50-x).....(1.40-x).......x

Kc = "\frac{[PCl_5]}{[PCl_3][Cl_2]}" = "\frac{x}{(2.50-x)(1.40-x)}" = 4.30×10-6

The equation

0.0000043x2 - 1.00001677x + 0.00001505 = 0

has given 2 roots x = 0.0206 and x = 22,325.18 ( x = 22,325.18 which is not acceptable because its gives negative concentration for these) Now we taken x = 0.0206

[PCl3] = 2.50 - x = 2.50 - 0.0206 = 2.294M

[Cl2] = 1.40 - x = 1.40 - 0.0206 = 1.194M

[PCl5] = x = 0.206M